OpenStudy (anonymous):

find the equation of a line having the giving slope containing the given point m=2/3 (8,-1)

OpenStudy (aroub):

y=mx+b , m=2/3 , (8,-1) -1=2/3(8)+b -1=16/3+b -16/3-1=b -19/3=b => y=2/3x-19/3

OpenStudy (anonymous):

thank you